In an ideal 3:17 step-up transformer, the primary power is 34 kW and the secondary current is 30 A. The primary voltage is:


A. 6.25 V

B. 200 V

C. 1133.3 V

D. 6422.2 V

E. 51 V
1
Expert's answer
2018-04-16T09:40:12-0400
I_pri=30/17 3= 5.294 A
V_pri=34000/5.294= 6422.2 V